# Guiding Masked Representation Learning to Capture Spatio-Temporal Relationship of Electrocardiogram

### 02 Feb 2024 v1

@@ Abstract
Electrocardiograms (ECG) are widely employed as a diagnostic tool for monitoring electrical signals originating from a heart. Recent machine learning research efforts have focused on the application of screening various diseases using ECG signals. However, adapting to the application of screening disease is challenging in that labeled ECG data are limited. Achieving general representation through self-supervised learning (SSL) is a well-known approach to overcome the scarcity of labeled data; however, a naive application of SSL to ECG data, without considering the spatial-temporal relationships inherent in ECG signals, may yield suboptimal results. In this paper, we introduce ST-MEM (Spatio-Temporal Masked Electrocardiogram Modeling), designed to learn spatio-temporal features by reconstructing masked 12-lead ECG data. ST-MEM outperforms other SSL baseline methods in various experimental settings for arrhythmia classification tasks. Moreover, we demonstrate that ST-MEM is adaptable to various lead combinations. Through quantitative and qualitative analysis, we show a spatio-temporal relationship within ECG data.
@@ 1 Introduction
The electrocardiogram (ECG) is a non-invasive heart measurement to monitor the electrical activity over time and diagnose diseases. Several supervised learning models have been developed to detect various heart diseases through ECG (Siontis et al., 2021). However, since the type of heart disease is diverse and the experienced cardiologists who can provide labels are limited, learning the ECG representation for each application (i.e., detecting various heart diseases) is challenging. Recently, self-supervised learning (SSL) for general representation has emerged in natural language processing (NLP) (Devlin et al., 2018; Brown et al., 2020) and computer vision (CV) (Chen et al., 2020; Grill et al., 2020; He et al., 2020; Caron et al., 2021) since it can be leveraged for numerous tasks, such as translation, sentence classification, image classification, and image generation. In ECG-based diagnosis, there were also similar efforts to learn general representation through SSL to overcome the limited resources and detect various heart diseases.
ECG-based representation learning through SSL is usually considered in two different learning methods, contrastive and generative learning (Jing & Tian, 2020). Contrastive learning (Sarkar & Etemad, 2020; Le et al., 2023; Gopal et al., 2021; Soltanieh et al., 2022; Kiyasseh et al., 2021; Wei et al., 2022) is a method to make the context before and after data augmentation to be similar. These data are usually defined by numerous augmentation methods (e.g., cropping, flipping, and shifting); however, despite a simple augmentation, the information on the semantic meaning of ECG signals can change drastically (Lan et al., 2023). Generative learning is a method of learning the presentation of data by reconstructing all or part of the input, such as a simple framework for masked image modeling (SimMIM) (Xie et al., 2022) or Masked Autoencoder (MAE) (He et al., 2022). In the case of ECG, there are variants of MAE which reconstructs input ECG signals (Zhang et al., 2022; Sawano et al., 2022; Hu et al., 2023). Generative learning is often augmentation-free or involves simpler data transformations that are more suitable for preserving the integrity of ECG data. This approach ensures that the context and meaningful information in ECG signals are retained.
@@
In ECG-based representation learning, exploiting both spatial and temporal information in ECG is significant. For instance, if we have a $L$-lead ECG, it means that cardiac activity is obtained over a duration with $L$ views. Therefore, we can understand a heart complementary when its ECG is gained not only spatially or temporally but spatio-temporally. The most common setting is standard 12-lead ECG (i.e., a heart is observed in 12 views) as shown in Figure 1, and for some cases, ECG in which its leads are a subset of 12-lead (i.e., reduced lead ECG) is acquired.
In this work, we leverage ECG to learn general representations by introducing a simple but effective self-supervised learning framework using MAE architecture. Throughout this process, both temporal and spatial information present in the ECG is utilized. The approach involves applying spatio-temporal patchifying to ECG data, as illustrated in Figure 2 (c), with the lead indicator such as lead-wise shared decoder, learnable lead embeddings, and separation embedding, as depicted in Figure 3. Moreover, we show that our model can be structurally fine-tuned for reduced lead ECG, demonstrating excellent performance not only in the standard 12-lead setting but also in limb leads and single leads. Finally, through quantitative and qualitative analysis, we demonstrated that spatial and temporal features were effectively learned.
Our contributions are the following. (1) We propose the simple but effective ECG specific generative self-supervised learning framework, named ST-MEM (Spatio-Temporal Masked Electrocardiogram Modeling). (2) ST-MEM can learn general representation by capturing spatio-temporal relationship of ECGs. We show this relationship by quantitative and qualitative analysis. (3) Through extensive experiments, ST-MEM demonstrates comparable performance to other contrastive and generative learning methods, which are widely used in ECG representation learning. We validate that ST-MEM excels in both fine-tuning and linear evaluation for the arrhythmia classification task. This robust performance extends across various scenarios, such as scarcity of labeled data and reduced lead settings.
@@ 2 Backgrounds: ECG
Although the standard 12-lead ECG remains significant, the utilization of its reduced lead sets should be considered as well. The advancement of mobile devices, such as smart watches that is capable of ECG measurements, has led to a substantial increase in limb leads and single lead data. Obtaining reduced lead sets like limb leads or a single lead is often preferred. Therefore, in this work, we demonstrate our ECG representation not only on the 12-lead setting but also on reduced lead settings.
@@ 3.1 Self-supervised pre-training with spatio-temporal masked auto-encoder
In this section, we present ST-MEM, a self-supervision framework for general representation learning for ECG. This incorporates the spatio-temporal relationship which leads to providing enhanced representation. An overview of our proposed method is depicted in Figure 3.
Masked auto-encoder with spatio-temporal patchifying. We consider a pre-text task to reconstruct a randomly masked portion of the data. In particular, we adopt an auto-encoder-based reconstruction task, called MAE (He et al., 2022). It consists of a vision transformer (ViT) encoder (Dosovitskiy et al., 2020) and a decoder with additional transformer blocks. The details of implementation can be found in Appendix E.
Considering an ECG signal denoted as ${\bm{X}}\in\mathbb{R}^{L\times T}$ with $L$ leads and length $T$, each lead signal, which is represented as ${\bm{X}}^{l}$, is divided into non-overlapping patches defined as $Patch^{l}=\{Patch_{1}^{l},...,Patch_{n}^{l}\}\in\mathbb{R}^{n\times p}$. Here, $n$ is determined by $T/p$, and $p$ represents the size of each patch. We independently split each lead signal into spatio-temporal patches, as depicted in Figure 2 (c), resulting in a total of $L\times n$ patches for a single 12-lead ECG signal.
@@ 4.1 Datasets
There are three 12-lead ECG datasets designated for pre-training, along with three another ECG datasets intended for downstream tasks. In the context of pre-training, we leverage the combined power of all three datasets. These datasets consist of Chapman (Zheng et al., 2020b), Ningbo (Zheng et al., 2020a), and CODE-15 (Ribeiro et al., 2021). As for the downstream datasets, PTB-XL (Wagner et al., 2020) and CPSC2018 (Liu et al., 2018) are 12-lead ECG datasets. In contrast, PhysioNet2017 (Clifford et al., 2017) is a single-lead dataset of lead I. For consistency, we resampled all ECGs, including pre-training and downstream datasets, to 250Hz.
Pre-training datasets. Chapman dataset comprises 10,646 12-lead ECG recordings, each lasting 10 seconds with a sample rate of 500 Hz. Similarly, Ningbo dataset contains 34,905 12-lead ECG recordings, also 10 seconds in duration and sampled at 500Hz. On the other hand, CODE-15 encompasses 345,779 12-lead ECG recordings from 233,770 patients, with a sample rate of 400 Hz. Within the 345,779 ECGs, 143,328 are 10-second recordings, while the remainder are 7 seconds in duration. In alignment with the PhysioNet 2021 challenge (Reyna et al., 2021), we removed 399 noisy ECGs from the Chapman dataset. We proceeded to merge these three datasets into a unified dataset for pre-training. This pre-training dataset comprises a total of 188,480 ECGs. Note that, during pre-training, we use all ECGs without focusing on specific labels.
Downstream datasets. PTB-XL comprises 21,837 12-lead ECG recordings collected from 18,885 patients, each lasting 10 seconds and sampled at a rate of 500 Hz. This dataset has five distinct labels, including cardiac arrhythmia and myocardial infarction. In the case of CPSC2018, it contains 6,877 12-lead ECG recordings, ranging from 6 seconds to 60 seconds in duration, with a sample rate of 500 Hz. This dataset is characterized by nine different labels associated with cardiac arrhythmia. Lastly, PhysioNet2017 consists of 8,528 ECGs recorded from lead I, with durations spanning from 9 to 60 seconds and sampled at a rate of 200 Hz. This dataset has four cardiac arrhythmia categories. ECGs shorter than 10 seconds are omitted, and for the remaining records, each are cropped into a consistent length of disjoint 10 seconds. These cropped ECG segments are considered as individual data points and evaluated independently. Detailed information on the specific number of instances used during self-supervised pre-training and supervised training can be found in the appendix A.
@@ 4.2 Baselines
All backbones are fixed as ViT-B (Dosovitskiy et al., 2020) with different patch projection layer corresponding to the patchifying way. We compared our pre-training method to networks initialized randomly (Supervised) and pre-trained models which learns representation by contrastive learning and generative learning method. In the realm of contrastive learning, we evaluated both MoCo v3 (He et al., 2020) and Contrastive Multi-segment Coding (CMSC) from CLOCS (Kiyasseh et al., 2021). MoCo v3 encourages the similarity between representations of instances and their augmented counterparts, while CMSC divides an ECG into two temporal segments and encourages the similarity of representations for these segments.
Additionally, we examined generative learning methods, specifically Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE) from MaeFE (Zhang et al., 2022). MTAE constructs representations by reconstructing temporal patches as depicted in Figure 2 (a), while MLAE does by reconstructing spatial patches, illustrated in Figure 2 (b). These models do not have lead indicating modules (e.g., lead-wise decoder, SEP and lead embeddings) and differ in their approach to patchifying input ECGs. Furthermore, to enhance robustness and adaptation to reduced lead sets, we introduce augmentation Random Lead Masking (RLM) (Oh et al., 2022) to MTAE, resulting in MTAE+RLM. All details for pre-train models can be found in Appendix E.
@@ 5.1 Experimental results
As shown in Table 1, we evaluate the general ECG representation obtained from each SSL method by conducting both linear evaluation and fine-tuning experiments on unseen datasets (i.e., not used during a pre-training stage). PTB-XL is a task of classifying both myocardial infarction (MI) and cardiac arrhythmia, while CPSC2018 consists of only cardiac arrhythmia Cardiac arrhythmia can be divided into different sub-classes (e.g., AV block (AVB) and premature atrial contraction (PAC) in both PTB-XL and CPSC2018 datasets.. Our proposed method, ST-MEM, shows outperforming performance against other baseline methods. In particular, we achieve a non-trivial margin with others on accuracy, F1, and AUROC scores in the linear evaluation. For instance, the F1 score is increased in the range of 0.064 to 0.366 and 0.292 to 0.561 on both PTB-XL and CPSC2018 datasets. This demonstrates that ST-MEM learns general ECG representation compared to others by explicitly considering both spatial and temporal information for ECG signals. Moreover, we present the applicability of ST-MEM by adapting to each different heart disease task. Although the baseline methods are pre-trained from large ECG datasets, their performance is similar to the supervised learning method, only used with downstream datasets, PTB-XL and CPSC2018. The CMSC pre-training method is even lower than the supervised learning method when we fine-tune the encoder. However, ST-MEM shows consistent outperforming performance on task adaptation, which is fine-tuning on downstream datasets.
@@ 5.5 Ablation study
Table 5 shows the importance of the lead indication module for ST-MEM. We conduct the linear evaluation in both PTB-XL and CPSC2018 datasets and compute the AUROC score. Applying only spatio-temporal patchifying without considering the lead type yields the lowest performance. However, when we include lead-wise shared encoder and each lead indication module (e.g., SEP embedding, and lead embedding), the results of linear evaluation monotonically increase in a CPSC2018 dataset. This demonstrates the effectiveness of spatial information while learning the general ECG representation. In Appendix B, we also visualize the embedding space from both ST-MEM and the model that only applies the spatio-temporal patchifying. The visualization is depicted in Figure 6. The additional ablation study can be found in Appendix C.
@@ 6 Related Works
6.1 ECG on self-supervised learning
Recently, research employing deep learning for ECG analysis has extended to a wide range of tasks, including the diagnosis of cardiac diseases such as arrhythmia classification (Ribeiro et al., 2020; Hu et al., 2022; Jun et al., 2018; Strodthoff et al., 2020), emotion recognition (Sarkar & Etemad, 2020), and patient identification (Oh et al., 2022; Li et al., 2020). Furthermore, self-supervised learning is utilized to learn general representations, with some studies focusing on contrastive learning methods. In contrastive learning, there are papers that apply concepts such as semantic preservation after data augmentation,  (Chen et al., 2020; He et al., 2020), to ECG data (Lai et al., 2023). Additionally, there is a study on the utilization of temporal invariance and spatial invariance within a single ECG data to preserve semantics (Kiyasseh et al., 2021). Some research combines CNN and transformer architectures to learn local and global features from ECG data (Oh et al., 2022). In terms of generative learning, MaeFE (Zhang et al., 2022) applies MAE (He et al., 2022), which patchifies temporally or spatially for analysis of ECG. Moreover, some studies (Sawano et al., 2022) were conducted by applying spatio-temporal patchifying, shown in Figure 2 (c), to acquire general ECG representation, but without lead indicating modules, it is challenging to learn spatial relationship between leads.
@@ 6.1 ECG on self-supervised learning
6.2 Challenging on reduced lead ECG
Obtaining a standard 12-lead ECG, however, can be excessive and often requires high-level clinical knowledge (Giannetta et al., 2020) that may not be readily available. Therefore, in recent times, advancements in ECG technologies have resulted in the creation of smaller and portable devices that can obtain the reduced lead set of 12-lead ECG. This advancement promotes the research to apply a machine learning technique to a reduced lead ECG setting (Kiyasseh et al., 2021; Hannun et al., 2019; Urtnasan et al., 2022; Mathews et al., 2018). It is worth noting that the reduced-lead ECG dataset is less abundant compared to 12-lead ECG, and there is research dedicated to addressing this limitation. In the context of self-supervised learning, RLM (Oh et al., 2022) was introduced to acquire robust representation to input ECG lead combinations. On the other hand, in Knowledge Distillation approaches, there is a work that utilizes a model pre-trained on 12-lead ECG as a teacher model and introduces a student model that takes a single lead as input (Qin et al., 2023). This student model leverages the representation from the 12-lead ECG teacher model.

### 11 Mar 2024 v2

@@ Abstract
Electrocardiograms (ECG) are widely employed as a diagnostic tool for monitoring electrical signals originating from a heart. Recent machine learning research efforts have focused on the application of screening various diseases using ECG signals. However, adapting to the application of screening disease is challenging in that labeled ECG data are limited. Achieving general representation through self-supervised learning (SSL) is a well-known approach to overcome the scarcity of labeled data; however, a naive application of SSL to ECG data, without considering the spatial-temporal relationships inherent in ECG signals, may yield suboptimal results. In this paper, we introduce ST-MEM (Spatio-Temporal Masked Electrocardiogram Modeling), designed to learn spatio-temporal features by reconstructing masked 12-lead ECG data. ST-MEM outperforms other SSL baseline methods in various experimental settings for arrhythmia classification tasks. Moreover, we demonstrate that ST-MEM is adaptable to various lead combinations. Through quantitative and qualitative analysis, we show a spatio-temporal relationship within ECG data. Our code is available at https://github.com/bakqui/ST-MEM.
@@ 1 Introduction
The electrocardiogram (ECG) is a non-invasive heart measurement to monitor the electrical activity over time and diagnose diseases. Several supervised learning models have been developed to detect various heart diseases through ECG (Siontis et al., 2021). However, since the types of heart disease are diverse and the experienced cardiologists who can provide labels are limited, learning the ECG representation for each application (i.e., detecting various heart diseases) is challenging. Recently, self-supervised learning (SSL) for general representation has emerged in natural language processing (kenton2019bert; Brown et al., 2020) and computer vision (Chen et al., 2020; Grill et al., 2020; He et al., 2020; Caron et al., 2021) since it can be leveraged for numerous tasks, such as translation, sentence classification, image classification, and image generation. In ECG-based diagnosis, there were also similar efforts to learn general representation through SSL to overcome the limited resources and detect various heart diseases.
ECG-based representation learning through SSL is usually considered in two different learning methods: contrastive and generative learning (Jing & Tian, 2020). Contrastive learning (Sarkar & Etemad, 2020; Le et al., 2023; Gopal et al., 2021; Soltanieh et al., 2022; Kiyasseh et al., 2021; Wei et al., 2022) is a method to ensure similarity in the context before and after data augmentation. These data are usually defined by numerous augmentation methods (e.g., cropping, flipping, and shifting); however, despite a simple augmentation, the information on the semantic meaning of ECG signals can change drastically (Lan et al., 2023). Generative learning is a method of learning the representation of data by reconstructing all or part of the input, such as a simple framework for masked image modeling (Xie et al., 2022) or Masked Autoencoder (MAE) (He et al., 2022). In the case of ECG, there are variants of MAE that reconstruct input ECG signals (Zhang et al., 2022; Sawano et al., 2022; Hu et al., 2023). Generative learning is often augmentation-free or involves simpler data transformations that are more suitable for preserving the integrity of ECG data. This approach ensures that the context and meaningful information in ECG signals are retained.
@@
In ECG-based representation learning, exploiting both spatial and temporal information in ECG is significant. For instance, if we have an $L$-lead ECG, it means that cardiac activity is obtained over a duration with $L$ views. Therefore, we can understand a heart complementary when its ECG is gained not only spatially or temporally but spatio-temporally. The most common setting is standard 12-lead ECG (i.e., a heart is observed in 12 views) as shown in Figure 1, and for some cases, ECG in which its leads are a subset of 12-lead (i.e., reduced lead ECG) is acquired.
In this work, we leverage ECG to learn general representations by introducing a simple but effective self-supervised learning framework using MAE architecture. Throughout this process, both temporal and spatial information present in the ECG is utilized. The approach involves applying spatio-temporal patchifying to ECG data, as illustrated in Figure 2 (c), with lead indicators such as lead-wise shared decoder, learnable lead embeddings, and separation embedding, as depicted in Figure 3. Moreover, we show that our model can be structurally fine-tuned for reduced lead ECG, demonstrating excellent performance not only in the standard 12-lead setting but also in limb leads and single leads. Finally, through quantitative and qualitative analysis, we demonstrated that spatial and temporal features were effectively learned.
Our contributions are the following. (1) We propose the simple but effective ECG-specific generative self-supervised learning framework, named ST-MEM (Spatio-Temporal Masked Electrocardiogram Modeling). (2) ST-MEM can learn general representation by capturing spatio-temporal relationship of ECGs. We show this relationship by quantitative and qualitative analysis. (3) Through extensive experiments, ST-MEM demonstrates comparable performance to other contrastive and generative learning methods, which are widely used in ECG representation learning. We validate that ST-MEM excels in both fine-tuning and linear evaluation for the arrhythmia classification task. This robust performance extends across various scenarios, such as scarcity of labeled data and reduced lead settings.
@@ 2 Backgrounds: ECG
Although the standard 12-lead ECG remains significant, the utilization of its reduced lead sets should be considered as well. The advancement of mobile devices, such as smartwatches that are capable of ECG measurements, has led to a substantial increase in limb leads and single lead data. Obtaining reduced lead sets like limb leads or a single lead is often preferred. Therefore, in this work, we demonstrate our ECG representation not only on the 12-lead setting but also on reduced lead settings.
@@ 3.1 Self-supervised pre-training with spatio-temporal masked auto-encoder
In this section, we present ST-MEM, a self-supervision framework for general representation learning for ECG. This incorporates the spatio-temporal relationship, which leads to providing enhanced representation. An overview of our proposed method is depicted in Figure 3.
Masked auto-encoder with spatio-temporal patchifying. We consider a pre-text task to reconstruct a randomly masked portion of the data. In particular, we adopt an auto-encoder-based reconstruction task, called MAE (He et al., 2022). It consists of a vision transformer (ViT) encoder (Dosovitskiy et al., 2020) and a decoder with additional transformer blocks.
Considering an ECG signal denoted as ${\bm{X}}\in\mathbb{R}^{L\times T}$ with $L$ leads and length $T$, each $l$th lead signal, represented as ${\bm{X}}^{l}$, is divided into non-overlapping patches defined as $Patch^{l}=\{Patch_{1}^{l},...,Patch_{n}^{l}\}\in\mathbb{R}^{n\times p}$. Here, $n$ is determined by $T/p$, and $p$ represents the size of each patch. We independently split each lead signal into spatio-temporal patches, as depicted in Figure 2 (c), resulting in a total of $L\times n$ patches for a single 12-lead ECG signal.
@@ 4 Experimental settings
This section provides a brief overview of the datasets utilized in our experiments, as well as the baselines against which our proposed method will be compared. More detailed information regarding the datasets, baselines, and implementation can be found in Appendix A.
@@ 4.1 Datasets
There are three 12-lead ECG datasets designated for pre-training, along with three other ECG datasets intended for downstream tasks. In the context of pre-training, we leverage the combined power of all three datasets. These datasets consist of Chapman (Zheng et al., 2020b), Ningbo (Zheng et al., 2020a), and CODE-15 (Ribeiro et al., 2021). As for the downstream datasets, we used PTB-XL (Wagner et al., 2020), CPSC2018 (Liu et al., 2018), and PhysioNet2017 (Clifford et al., 2017) datasets, where PhysioNet2017 is a single-lead dataset (i.e., lead I dataset) and others are 12-lead ECG datasets. For consistency, we resampled all ECGs, including pre-training and downstream datasets, to 250Hz.
Pre-training datasets. Chapman dataset comprises 10,646 12-lead ECG recordings, each lasting 10 seconds with a sample rate of 500 Hz. Similarly, the Ningbo dataset contains 34,905 12-lead ECG recordings, also 10 seconds in duration and sampled at 500Hz. On the other hand, CODE-15 encompasses 345,779 12-lead ECG recordings from 233,770 patients, with a sample rate of 400 Hz. Within the 345,779 ECGs, 143,328 are 10-second recordings, with the remainder being 7 seconds, and we and we exclusively utilized the 10-second ECGs. In alignment with the PhysioNet 2021 challenge (Reyna et al., 2021), we removed 399 ECGs from the Chapman dataset. We proceeded to merge these three datasets into a unified dataset for pre-training. This pre-training dataset comprises a total of 188,480 ECGs. Note that, during pre-training, we use all ECGs without focusing on specific labels.
Downstream datasets. PTB-XL comprises 21,837 12-lead ECG recordings collected from 18,885 patients, each lasting 10 seconds and sampled at a rate of 500 Hz. This dataset has five distinct labels, including cardiac arrhythmia and myocardial infarction. In the case of CPSC2018, it contains 6,877 12-lead ECG recordings, ranging from 6 seconds to 60 seconds in duration, with a sample rate of 500 Hz. This dataset is characterized by nine different labels associated with cardiac arrhythmia. Lastly, PhysioNet2017 consists of 8,528 ECGs recorded from lead I, with durations spanning from 9 to 60 seconds and sampled at a rate of 200 Hz. This dataset has four cardiac arrhythmia categories. ECGs shorter than 10 seconds are omitted, and for the remaining records, each is cropped into a consistent length of disjointed 10 seconds. These cropped ECG segments are considered individual data points and evaluated independently.
@@ 4.2 Baselines
All backbones are fixed as ViT-B (Dosovitskiy et al., 2020) with different patch projection layers corresponding to the patchifying way. We compared our pre-training method to networks initialized randomly (i.e., Supervised) and pre-trained models that learn representation by contrastive learning and generative learning methods. In the realm of contrastive learning, we evaluated both MoCo v3 (chen2021empirical) and Contrastive Multi-segment Coding (CMSC) from CLOCS (Kiyasseh et al., 2021). MoCo v3 encourages the similarity between representations of instances and their augmented counterparts, while CMSC divides an ECG into two temporal segments and encourages the similarity of representations for these segments.
Additionally, we examined generative learning methods, specifically Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE) from MaeFE (Zhang et al., 2022). MTAE constructs representations by reconstructing temporal patches as depicted in Figure 2 (a), while MLAE does by reconstructing spatial patches, illustrated in Figure 2 (b). These models do not have lead-indicating modules (e.g., lead-wise decoder, SEP, and lead embeddings) and differ in their approach to patchifying input ECGs. Furthermore, to enhance robustness and adaptation to reduced lead sets, we introduce augmentation Random Lead Masking (RLM) (Oh et al., 2022) to MTAE, resulting in MTAE+RLM.
@@ 5 Experiments and results
In this section, we examine the results of our experiments, evaluating them both quantitatively and qualitatively to verify the effectiveness of ST-MEM. Additional experimental results are reported in Appendix B.
@@ 5.1 Experimental results
As shown in Table 1, we evaluate the general ECG representation obtained from each SSL method by conducting both linear evaluation and fine-tuning experiments on unseen datasets (i.e., not used during a pre-training stage). PTB-XL is a task of classifying both myocardial infarction (MI) and cardiac arrhythmia, while CPSC2018 consists of only cardiac arrhythmia  [^1] Cardiac arrhythmia can be divided into different sub-classes (e.g., AV block (AVB) and premature atrial contraction (PAC)) in both PTB-XL and CPSC2018 datasets.. Our proposed method, ST-MEM, shows outperforming performance against other baseline methods. In particular, we achieve a non-trivial margin with others on accuracy, F1, and AUROC scores in the linear evaluation. For instance, the F1 score is increased in the range of 0.064 to 0.366 and 0.292 to 0.561 on both PTB-XL and CPSC2018 datasets. This demonstrates that ST-MEM learns general ECG representation compared to others by explicitly considering both spatial and temporal information for ECG signals. Moreover, we present the applicability of ST-MEM by adapting to each different heart disease task. Although the baseline methods are pre-trained from large ECG datasets, their performance is similar to the supervised learning method, only used with downstream datasets, PTB-XL and CPSC2018. The CMSC pre-training method is even lower than the supervised learning method when we fine-tune the encoder. However, ST-MEM shows consistent outperforming performance on task adaptation, which is fine-tuning on downstream datasets.
@@ 5.5 Ablation study
Table 5 shows the importance of the lead indication module for ST-MEM. We conduct the linear evaluation in both PTB-XL and CPSC2018 datasets and compute the AUROC score. Applying only spatio-temporal patchifying without considering the lead type yields the lowest performance. However, when we include lead-wise shared encoder and each lead indication module (e.g., SEP embedding, and lead embedding), the results of linear evaluation monotonically increase in a CPSC2018 dataset. This demonstrates the effectiveness of spatial information while learning the general ECG representation.
@@ 6 Related Works
ECG on self-supervised learning. Recently, research employing de ep learning for ECG analysis has extended to a wide range of tasks, including the diagnosis of cardiac diseases such as arrhythmia classification (Ribeiro et al., 2020; Hu et al., 2022; Jun et al., 2018; Strodthoff et al., 2020), emotion recognition (Sarkar & Etemad, 2020), and patient identification (Oh et al., 2022; Li et al., 2020). Furthermore, self-supervised learning is utilized to learn general representations, with some studies focusing on contrastive learning methods. In contrastive learning, some papers apply concepts such as semantic preservation after data augmentation (Chen et al., 2020; He et al., 2020), to ECG data (Lai et al., 2023). Additionally, there is a study on the utilization of temporal invariance and spatial invariance within a single ECG data to preserve semantics (Kiyasseh et al., 2021). Some research combines CNN and transformer architectures to learn local and global features from ECG data (Oh et al., 2022). In terms of generative learning, MaeFE (Zhang et al., 2022) applies MAE (He et al., 2022), which patchifies temporally or spatially for analysis of ECG. Moreover, some studies (Sawano et al., 2022) were conducted by applying spatio-temporal patchifying, shown in Figure 2 (c), to acquire general ECG representation, but without lead-indicating modules, it is challenging to learn the spatial relationship between leads.
@@
Challenging on reduced lead ECG. Obtaining a standard 12-lead ECG, however, can be excessive and often requires high-level clinical knowledge (Giannetta et al., 2020) that may not be readily available. Therefore, recent advancements in ECG technologies have resulted in the creation of smaller and portable devices that can obtain the reduced lead set of 12-lead ECG. This advancement promotes the research to apply a machine learning technique to a reduced lead ECG setting (Kiyasseh et al., 2021; Hannun et al., 2019; Urtnasan et al., 2022; Mathews et al., 2018). It is worth noting that the reduced-lead ECG dataset is less abundant compared to 12-lead ECG, and there is research dedicated to addressing this limitation. In the context of self-supervised learning, RLM (Oh et al., 2022) was introduced to acquire robust representation to input ECG lead combinations. On the other hand, in Knowledge Distillation approaches, there is a work that utilizes a model pre-trained on 12-lead ECG as a teacher model and introduces a student model that takes a single lead as input (Qin et al., 2023). This student model leverages the representation from the 12-lead ECG teacher model.